# Quasiparticle states of hexagonal BN: A van der Waals density functional study

### v1

@@ I Introduction
In practice, we need MBPT input to define robust so-called exchange-correlation (XC) functional approximations [mabr, la70, langreth1975, gulu76, langreth1977, lape80, lameprl1981, lavo87, langreth1990, perdew21996, casida1999, thonhauser2007, hyldgaard2020]. Approximations also enter MBPT-based QP studies (e.g., in GW [reining2018]) and we seek consistency in XC design work [helujpc1971, lape80, shsl1983, casida1999, chong2002, ferretti2014, hyldgaard2020, JPCM2025]. Progress in such a concerted approach should and can be tested on both GS and QP properties [cococcioni2005, nguyen2018, WiOhHa21, JPCM2025].
@@
With an accurate XC functional approximation, KS-DFT lets us predict the atomic configurations [burke] and materials vibrations, e.g., Refs. [BrownAltvPRB16, grwahy20]. Such computational characterizations permit theory validation both directly (on thermophysical properties and phases, e.g., Ref. [PeGrRo20]) and indirectly on the QPs, as revealed in electrical transport measurements [yu2010], scanning tunneling microscopy (STM) [magonov1993], angle-resolved photoemission (ARPES) [zhang2022, grunert2020], or inverse photoemission [smith1988]. These properties generally reflect the QP band structure (energy variation with momentum in an extended system). In terms of MBPT and XC quality testing, one key observation is that (sub-)picometer accuracy at atomic structure characterization is typically required to, in turn, accurately predict the light-matter interactions in the more costly MBPT methods [RanPRB16].
@@
Notably, direct theory-experiment testing on QPs has a higher value still for materials-theory development. We observe that Koopmans-integer (KI) DFT [borghi2014, nguyen2018, gennaro2022, colonna2022, linscott2023], using a linear-response description based on a given input XC functional, stands out by providing fast QP predictions without changing neither total energies nor the forces that set structure when that XC input is used in KS-DFT [linscott2023]. Accurate tools, linking the structure and QP aspects of materials modeling, are valuable for characterizations in their own rights but KI-DFT offers a potential for synenergy in continued method developments. The structure of the Hedin QP-equation-of-motion framework is equivalent to the formal adiabatic-connection formula (ACF) specification of the XC functional designs [helujpc1971, perdew21996, hyldgaard2020]. For a versatile XC functional, and certainly when used for layered materials, we expect to find fingerprints of, for example, van der Waals (vdW) interactions on the QPs band structure, that is, in the energy variation with the QP momentum. We can test, for example, the accuracy of an XC choice on descriptions of the interlayer vdW attractions by comparing the corresponding QP predictions by KI-DFT to either measurements or direct MBPT characterizations.
@@
Present-standard KI-DFT studies [Dabo2013, borghi2014, nguyen2015, nguyen2016, nguyen2018, colonna2018, colonna2019, Elliott2019, almeida2021, colonna2022, gennaro2022, linscott2023, marrazzo2024], rely on either the local-density approximation (LDA) [helujpc1971, pewa92] or the the semilocal Perdew-Burke-Ernzerhof (PBE) [perdew21996] XC functional. The latter of these functionals is developed for KS-DFT and gives what we here label as PBE results. Via an extension to density functional perturbation theory (DFPT), PBE also permits efficient QP characterizations in what we call KI-PBE, when a piecewise-linearity (PWL) constraint [perdew1982, davo2010, nguyen2018, linscott2023, schubert2023] is imposed on the energy variation with partial charging. However, KI-DFT still reflects the nature of the XC functional that is chosen for KI-DFT [linscott2023] and KI-PBE is not set up to track vdW interactions [hyldgaard2020]. This holds also for KI-PBE predictions of the QP band structure characterization, that is, in the QP energy variation with momentum.
@@
It is motivated to move KI-DFT usage to instead start from a truly nonlocal-correlation functional form, available within the vdW density functional (vdW-DF) method [rydberg2003, dion2004, thonhauser2007, hybesc14, Berland_2015:van_waals, Thonhauser_2015:spin_signature, hyldgaard2020, shukla2022, JPCM2025]. Stacking multiple layers of a two-dimensional material to form a bulk system is known to alter its band structure [wickramaratne2018]. An expansion that reflects an incorrect structure prediction of the interlayer spacing, will induce further changes [hunt2020]. Beyond those observations is again the fact that QPs are the most direct way that we can test the quality of the QP and MBPT logic behind many of our XC functionals designs [helujpc1971, shsl1983, casida1999, perdew21996, rydberg2000, dion2004, hybesc14, berland2014, hyldgaard2020].
@@
Here we therefore present a study tracking vdW signatures on QPs in hexagonal boron-nitride systems (h-BN) while introducing the use of KI-CX, that is, KI-DFT with a start in the consistent–exchange vdW-DF-cx version (here abbreviated CX) [berland2014, berland2014a, hyldgaard2020]. We note that the QuantumESPRESSO (QE) suite of DFT codes [giannozzi2009, giannozzi2017], with the module for Koopmans-DFT extension [koopweb2025, linscott2023], is already set up for KI-CX studies. The h-BN systems are insulating and have been thoroughly studied both theoretically and experimentally. h-BN has emerged as one of the most prominent two-dimensional materials, owing in part to a graphene-like honeycomb structure in its layers and to its wide band gap [caldwell2019, roy2021]. Additionally, h-BN exhibits a suite of advantageous properties, such as excellent thermal conductivity [lindsay2011], unique thickness-dependent behavior [li2009], chemical stability [weng2016], optical functionality [zunger1976], and polymer compatibility [yu2018]. These attributes enable applications in aerospace [zhang2017, weng2016], environmental science, and energy technologies—spanning adsorption, photo-catalysis, membrane separation, disinfection, environmental sensing, as well as energy conversion and storage [li2022].
@@
The QPs correspond to sudden events that may be either adding or removing an electron [AuJoWi00]. We focus on QPs in three systems that are experimentally available, and we first set the atomic structure by observations. This means that KI-CX versus KI-PBE differences reflect the direct impacts of truly nonlocal-correlation effects (interlayer coupling) on the QPs. We next track the full impact of including nonlocal-correlations (via KI-CX) on the QP description. This is done by also contrasting KI-CX and KI-PBE QP descriptions at the structure predictions provided by these XC functional choices (namely, using CX and PBE in KS-DFT, respectively). There are local and global valence-band (conduction-band) extrema, together forming a number of (indirect) QP gaps in h-BN systems. We assess KI-PBE and KI-CX quality on accuracy for those gaps, comparing with experimental observations and GW studies.
@@ II.1 Truly nonlocal correlation in CX
The design of the vdW-DFs relies on MBPT characterizations [mabr, gulu76, langreth1977, lape80, lameprl1981, lavo87, langreth1990, thonhauser2007] of the density-density response function $\chi_{\lambda}\quantity(\mathbf{r},\mathbf{r^{\prime}},\omega)$, i.e., the ratio of the density change $\delta n(\mathbf{r})$ that arise when external potential at frequency $\omega$ is applied at position $\mathbf{r^{\prime}}$. As indicated by the subscript, this response is determined while assuming that the raw inter-electron Coulomb matrix element $V=|\mathbf{r}-\mathbf{r^{\prime}}|^{-1}$ is scaled by a coupling constant $0<\lambda<1$. The adiabatic connection formula (ACF) [langreth1975, gulu76, langreth1977, gunnarsson1977],
@@
The resulting vdW-DF functionals are in practice given by treating the XC design work as an electrodynamics problem [rydberg2000, rydberg2003, hybesc14]. Critical are the focus on collective excitations, compliance with Lindhard screening, and enforcing current conservation and time-reversal symmetry in modeling the local-field response [dion2004, hybesc14, hyldgaard2020]. We start with a generalized-gradient approximation (GGA) for a so-called internal XC function $E_{\text{xc}}^{\text{in}}[\rho]$, comprising the XC terms of the local-density approximation (LDA) [helujpc1971, perdew1986] and a MBPT-guided choice for the gradient-corrected exchange [dion2004, lee10p081101, Berland_2015:van_waals, ChBeTh20], This ensures a GGA-type exchange component $E_{\text{x}}^{\text{in}}[\rho]$ in the starting point for the vdW-DF design. We next use an electrodynamics-based recast of the ACF [rydberg2000, rydberg2003, hyldgaard2020] to extract a corresponding initial-guess dielectric function $\epsilon(\mathbf{r},\mathbf{r^{\prime}},\omega=iu)$ [lee10p081101, Berland_2015:van_waals]:
@@
This reformulation of the semilocal functional $E^{\text{in}}_{\text{xc}}$ reflects an exponential resummation, $\epsilon(\omega)\equiv\exp(S_{\text{xc}}(\omega))$, of a plasmon-propagator $S_{\text{xc}}(\mathbf{r},\mathbf{r^{\prime}},\omega)$ that corresponds to the internal GGA [Berland_2015:van_waals], so that the initial-guess for the collective excitations (found at $\epsilon(\omega)=0$) are tracked by logarithmic singularities [jerry65, hybesc14]. Third, we note that the Coulomb Green function $G$ is set up to trace also long-ranged electrodynamics couplings of virtual electron dynamics (for example, as reflected by plasmons) [rydberg2003, hybesc14] and that the vdW forces are naturally included when we enforce current-conservation, $\epsilon(\omega)\longrightarrow\kappa(\omega)\equiv\nabla\epsilon(\omega)\cdot\nabla G$ [rydberg2000, rydberg2003, dion2004, hybesc14]. We therefore use the resulting beyond-GGA recast of the ACF for a formal vdW-DF-method specification [lee10p081101, Berland_2015:van_waals, hyldgaard2020]
@@
Finally, we often (but not always [rydberg2000, rydberg2003]) expand the truly nonlocal correlation term
@@
in terms of both the Lindhard screening an the exponential resummation to second order in $S_{\text{xc}}(\omega)$ [dion2004, Berland_2015:van_waals, hyldgaard2020]:
@@
The complete specification of the CX energy functional [berland2014, berland2014a] is denoted $E^{\text{CX}}_{\text{xc}}$. It has an exchange component that mirrors that of $E^{\text{in}}_{\text{xc}}$ (as far as possible) to ensure XC balance in calculations of energy differences [berland2014, hyldgaard2020]. CX yields a local effective potential that defines the equation of motion for the KS-DFT orbitals. The consistent focus on the plasmon behavior means that the specification of the total correlation, $E_{\text{c}}^{\text{LDA}}+E_{\text{c}}^{\text{nl}}$, is explicitly given by the density. Using Eq. (5) we can even cast the nonlocal-correlation energy functional [dion2004, thonhauser2007]
@@
with a spline representation [roso09] for what is a universal nonlocal-correlation kernel $\Phi_{c}^{\text{nl}}[\rho]$. This keeps computational costs comparable to that of any GGA-based KS-DFT [roso09], for example, PBE [perdew21996].
@@
In a set of recent designs we have sought to design and leverage generalized-KS vdW-DFs [levy1984, DFcx02017, cx0p2018, shukla_2022a, shukla2022, JPCM2025]. This is done in part to overcome residual accuracy issues in PBE or CX KS-DFT characterizations and sometimes also in the PBE-based range-separated hybrid (RSH), denoted HSE [HSE03, HSE06, OTRSHalga, skone2016, chen2018, WiOhHa21]; The relevance of having access to global-hybrid RSH vdW-DFs is illustrated and summarized, for example, in Refs. [PeGrRo20, grwahy20, frostenson2022, frostenson2024, JPCM2025]. Starting with two robust vdW-DF versions, CX and rev-vdW-DF2 [hamada14], and replacing, for example, part of the short-ranged GGA exchange, we arrive at two RSH vdW-DFs, denoted AHCX [shukla_2022a] and AHBR [shukla2022]. They generalize the HSE by providing a consistent inclusion of truly nonlocal-correlation effects and, in particular vdW interactions. The AHCX and AHBR are general-purpose options, noting that they are applicable also to metallic systems [shukla_2022a, shukla2022] and for a case when we expect strong static-correlation effects on molecular structure [frostenson2024]. Still, we can motivate adjustments of these basic RSH vdW-DF designs when used for molecular systems or for semiconductors: For the former (latter) we expect an absence (moderate presence) of asymptotic screening in the exchange descriptions, as implemented in our recent AHBR-mRSH [JPCM2025] (in global-hybrid vdW-DFs like CX0p [cx0p2018, grwahy20]). The logic is that the system-specific dielectric function is implicitly set by the Hartree-energy term and should therefore (ideally) also be reflected in the exchange description [gulu76, kronik2012, OTRSHgap, skone2016, chen2018, WiOhHa21, JPCM2025]. It is hard to realize this goal in basic XC approximations that aims to be fast and universally applicable [langreth1977, levy1984, perdew21996, berland2014].
@@
Importantly, practical use of such new hybrid vdW-DFs relies on the orbital-based evaluation of Fock exchange that are now available in QE [giannozzi2017, linACE, PaoloElStruct1, shukla_2022a]. That is, our hybrid vdW-DF calculations, e.g., [DFcx02017, PeGrRo20, shukla2022, frostenson2022, frostenson2024], automatically proceed in the generalized-KS framework [levy1984]. This opens for vdW-inclusive QP predictions, when we apply also the ideas so-called optimal tuning [kronik2012, OTRSHalga, OTRSHgap, OTRSHadsorp17, WiOhHa21] to our new molecule-focused AHBR-mRSH [JPCM2025]. Excitingly, use of AHBR-mRSH with optimal tuning set by consistency, constitute a generalized KS-DFT that complies with the PWL constraint [perdew1982, davo2010, OTRSHalga, WiOhHa21, JPCM2025] and has high accuracy for both binding energies and QPs of nitrogen-base systems [JPCM2025].
@@
The present focus on moving also KI-DFT to a vdW-DF start and our introduction of the optimally-tuned AHBR-mRSH supplement each other. The same applies also (to some extent) for the more traditional RSH vdW-DF designs, AHCX and AHBR, and when a Hubbard-U term (U) [liechtenstein1995, dudarev1998] is added to vdW-DF as the starting XC functional choice [pengperdew2017]. These methods restore PWL [cococcioni2005] to varying extents and, for example, RSH vdW-DFs simultaneously improve parameter-free vdW-DF descriptions of total energies and structure [shukla2022]. Also, AHBR-mRSH delivers QP predictions that employ the same nonlocal-correlation description [lee10p081101] that is already in wide KS-DFT usage [hamada14, shukla2022, JPCM2025]. However, by including some Fock exchange in AHBR-mRSH we also simultaneously correct some density-driven errors, e.g, for transition states, that affects the underlying rev-vdW-DF2 functional [shukla2022, JPCM2025]. In contrast, the here-introduced KI-CX approach (described below) leaves the GS energies and structure unchanged [colonna2022, linscott2023] and therefore tests CX directly in terms of both GS and QP properties.

### v2

@@ I Introduction
In practice, we need MBPT input to define robust so-called exchange-correlation (XC) functional approximations [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 34, 46, 35]. Approximations also enter MBPT-based QP studies (e.g., in GW [5]) and we seek consistency in XC design work [32, 41, 33, 34, 47, 25, 35, 23]. Progress in such a concerted approach should and can be tested on both GS and QP properties [48, 49, 21, 23].
@@
With an accurate XC functional approximation, KS-DFT lets us predict the atomic configurations [14] and materials vibrations, e.g., Refs. [50, 51]. Such computational characterizations permit theory validation both directly (on thermophysical properties and phases, e.g., Ref. [52]) and indirectly on the QPs, as revealed in electrical transport measurements [53], scanning tunneling microscopy (STM) [54], angle-resolved photoemission (ARPES) [55, 56], or inverse photoemission [57]. These properties generally reflect the QP band structure (energy variation with momentum in an extended system). In terms of MBPT and XC quality testing, one key observation is that (sub-)picometer accuracy at atomic structure characterization is typically required to, in turn, accurately predict the light-matter interactions in the more costly MBPT methods [3].
@@
Notably, direct theory-experiment testing on QPs has a higher value still for materials-theory development. We observe that Koopmans-integer (KI) DFT [26, 49, 27, 28, 29], using a linear-response description based on a given input XC functional, stands out by providing fast QP predictions without changing neither total energies nor the forces that set structure when that XC input is used in KS-DFT [29]. Accurate tools, linking the structure and QP aspects of materials modeling, are valuable for characterizations in their own rights but KI-DFT offers a potential for synenergy in continued method developments. The structure of the Hedin QP-equation-of-motion framework is equivalent to the formal adiabatic-connection formula (ACF) specification of the XC functional designs [32, 45, 35]. For a versatile XC functional, and certainly when used for layered materials, we expect to find fingerprints of, for example, van der Waals (vdW) interactions on the QPs band structure, that is, in the energy variation with the QP momentum. We can test, for example, the accuracy of an XC choice on descriptions of the interlayer vdW attractions by comparing the corresponding QP predictions by KI-DFT to either measurements or direct MBPT characterizations.
@@
Present-standard KI-DFT studies [58, 26, 59, 60, 49, 61, 62, 63, 64, 28, 27, 29, 65], rely on either the local-density approximation (LDA) [32, 66] or the the semilocal Perdew-Burke-Ernzerhof (PBE) [45] XC functional. The latter of these functionals is developed for KS-DFT and gives what we here label as PBE results. Via an extension to density functional perturbation theory (DFPT), PBE also permits efficient QP characterizations in what we call KI-PBE, when a piecewise-linearity (PWL) constraint [67, 24, 49, 29, 68] is imposed on the energy variation with partial charging. However, KI-DFT still reflects the nature of the XC functional that is chosen for KI-DFT [29] and KI-PBE is not set up to track vdW interactions [35]. This holds also for KI-PBE predictions of the QP band structure characterization, that is, in the QP energy variation with momentum.
@@
It is motivated to move KI-DFT usage to instead start from a truly nonlocal-correlation functional form, available within the vdW density functional (vdW-DF) method [69, 70, 46, 71, 72, 73, 35, 74, 23]. Stacking multiple layers of a two-dimensional material to form a bulk system is known to alter its band structure [75]. An expansion that reflects an incorrect structure prediction of the interlayer spacing, will induce further changes [76]. Beyond those observations is again the fact that QPs are the most direct way that we can test the quality of the QP and MBPT logic behind many of our XC functionals designs [32, 33, 34, 45, 77, 70, 71, 78, 35].
@@
Here we therefore present a study tracking vdW signatures on QPs in hexagonal boron-nitride systems (h-BN) while introducing the use of KI-CX, that is, KI-DFT with a start in the consistent–exchange vdW-DF-cx version (here abbreviated CX) [78, 79, 35]. We note that the QuantumESPRESSO (QE) suite of DFT codes [80, 81], with the module for Koopmans-DFT extension [30, 29], is already set up for KI-CX studies. The h-BN systems are insulating and have been thoroughly studied both theoretically and experimentally. h-BN has emerged as one of the most prominent two-dimensional materials, owing in part to a graphene-like honeycomb structure in its layers and to its wide band gap [82, 83]. Additionally, h-BN exhibits a suite of advantageous properties, such as excellent thermal conductivity [84], unique thickness-dependent behavior [85], chemical stability [86], optical functionality [87], and polymer compatibility [88]. These attributes enable applications in aerospace [89, 86], environmental science, and energy technologies—spanning adsorption, photo-catalysis, membrane separation, disinfection, environmental sensing, as well as energy conversion and storage [90].
@@
The QPs correspond to sudden events that may be either adding or removing an electron [1]. We focus on QPs in three systems that are experimentally available, and we first set the atomic structure by observations. This means that KI-CX versus KI-PBE differences reflect the direct impacts of truly nonlocal-correlation effects (interlayer coupling) on the QPs. We next track the full impact of including nonlocal-correlations (via KI-CX) on the QP description. This is done by also contrasting KI-CX and KI-PBE QP descriptions at the structure predictions provided by these XC functional choices (namely, using CX and PBE in KS-DFT, respectively). There are local and global valence-band (conduction-band) extrema, together forming a number of (indirect) QP gaps in h-BN systems. We assess KI-PBE and KI-CX quality on accuracy for those gaps, comparing with experimental observations and GW studies.
@@ II.1 Truly nonlocal correlation in CX
The design of the vdW-DFs relies on MBPT characterizations [36, 39, 40, 41, 42, 43, 44, 46] of the density-density response function $\chi_{\lambda}\quantity(\mathbf{r},\mathbf{r^{\prime}},\omega)$, i.e., the ratio of the density change $\delta n(\mathbf{r})$ that arise when external potential at frequency $\omega$ is applied at position $\mathbf{r^{\prime}}$. As indicated by the subscript, this response is determined while assuming that the raw inter-electron Coulomb matrix element $V=|\mathbf{r}-\mathbf{r^{\prime}}|^{-1}$ is scaled by a coupling constant $0<\lambda<1$. The adiabatic connection formula (ACF) [38, 39, 40, 91],
@@
The resulting vdW-DF functionals are in practice given by treating the XC design work as an electrodynamics problem [77, 69, 71]. Critical are the focus on collective excitations, compliance with Lindhard screening, and enforcing current conservation and time-reversal symmetry in modeling the local-field response [70, 71, 35]. We start with a generalized-gradient approximation (GGA) for a so-called internal XC function $E_{\text{xc}}^{\text{in}}[\rho]$, comprising the XC terms of the local-density approximation (LDA) [32, 92] and a MBPT-guided choice for the gradient-corrected exchange [70, 93, 72, 94], This ensures a GGA-type exchange component $E_{\text{x}}^{\text{in}}[\rho]$ in the starting point for the vdW-DF design. We next use an electrodynamics-based recast of the ACF [77, 69, 35] to extract a corresponding initial-guess dielectric function $\epsilon(\mathbf{r},\mathbf{r^{\prime}},\omega=iu)$ [93, 72]:
@@
This reformulation of the semilocal functional $E^{\text{in}}_{\text{xc}}$ reflects an exponential resummation, $\epsilon(\omega)\equiv\exp(S_{\text{xc}}(\omega))$, of a plasmon-propagator $S_{\text{xc}}(\mathbf{r},\mathbf{r^{\prime}},\omega)$ that corresponds to the internal GGA [72], so that the initial-guess for the collective excitations (found at $\epsilon(\omega)=0$) are tracked by logarithmic singularities [95, 71]. Third, we note that the Coulomb Green function $G$ is set up to trace also long-ranged electrodynamics couplings of virtual electron dynamics (for example, as reflected by plasmons) [69, 71] and that the vdW forces are naturally included when we enforce current-conservation, $\epsilon(\omega)\longrightarrow\kappa(\omega)\equiv\nabla\epsilon(\omega)\cdot\nabla G$ [77, 69, 70, 71]. We therefore use the resulting beyond-GGA recast of the ACF for a formal vdW-DF-method specification [93, 72, 35]
@@
Finally, we often (but not always [77, 69]) expand the truly nonlocal correlation term
@@
in terms of both the Lindhard screening an the exponential resummation to second order in $S_{\text{xc}}(\omega)$ [70, 72, 35]:
@@
The complete specification of the CX energy functional [78, 79] is denoted $E^{\text{CX}}_{\text{xc}}$. It has an exchange component that mirrors that of $E^{\text{in}}_{\text{xc}}$ (as far as possible) to ensure XC balance in calculations of energy differences [78, 35]. CX yields a local effective potential that defines the equation of motion for the KS-DFT orbitals. The consistent focus on the plasmon behavior means that the specification of the total correlation, $E_{\text{c}}^{\text{LDA}}+E_{\text{c}}^{\text{nl}}$, is explicitly given by the density. Using Eq. (5) we can even cast the nonlocal-correlation energy functional [70, 46]
@@
with a spline representation [96] for what is a universal nonlocal-correlation kernel $\Phi_{c}^{\text{nl}}[\rho]$. This keeps computational costs comparable to that of any GGA-based KS-DFT [96], for example, PBE [45].
@@
In a set of recent designs we have sought to design and leverage generalized-KS vdW-DFs [97, 98, 99, 22, 74, 23]. This is done in part to overcome residual accuracy issues in PBE or CX KS-DFT characterizations and sometimes also in the PBE-based range-separated hybrid (RSH), denoted HSE [100, 101, 18, 20, 19, 21]; The relevance of having access to global-hybrid RSH vdW-DFs is illustrated and summarized, for example, in Refs. [52, 51, 102, 103, 23]. Starting with two robust vdW-DF versions, CX and rev-vdW-DF2 [104], and replacing, for example, part of the short-ranged GGA exchange, we arrive at two RSH vdW-DFs, denoted AHCX [22] and AHBR [74]. They generalize the HSE by providing a consistent inclusion of truly nonlocal-correlation effects and, in particular vdW interactions. The AHCX and AHBR are general-purpose options, noting that they are applicable also to metallic systems [22, 74] and for a case when we expect strong static-correlation effects on molecular structure [103]. Still, we can motivate adjustments of these basic RSH vdW-DF designs when used for molecular systems or for semiconductors: For the former (latter) we expect an absence (moderate presence) of asymptotic screening in the exchange descriptions, as implemented in our recent AHBR-mRSH [23] (in global-hybrid vdW-DFs like CX0p [99, 51]). The logic is that the system-specific dielectric function is implicitly set by the Hartree-energy term and should therefore (ideally) also be reflected in the exchange description [39, 16, 105, 20, 19, 21, 23]. It is hard to realize this goal in basic XC approximations that aims to be fast and universally applicable [40, 97, 45, 78].
@@
Importantly, practical use of such new hybrid vdW-DFs relies on the orbital-based evaluation of Fock exchange that are now available in QE [81, 106, 107, 22]. That is, our hybrid vdW-DF calculations, e.g., [98, 52, 74, 102, 103], automatically proceed in the generalized-KS framework [97]. This opens for vdW-inclusive QP predictions, when we apply also the ideas so-called optimal tuning [16, 18, 105, 108, 21] to our new molecule-focused AHBR-mRSH [23]. Excitingly, use of AHBR-mRSH with optimal tuning set by consistency, constitute a generalized KS-DFT that complies with the PWL constraint [67, 24, 18, 21, 23] and has high accuracy for both binding energies and QPs of nitrogen-base systems [23].
@@
The present focus on moving also KI-DFT to a vdW-DF start and our introduction of the optimally-tuned AHBR-mRSH supplement each other. The same applies also (to some extent) for the more traditional RSH vdW-DF designs, AHCX and AHBR, and when a Hubbard-U term (U) [109, 110] is added to vdW-DF as the starting XC functional choice [111]. These methods restore PWL [48] to varying extents and, for example, RSH vdW-DFs simultaneously improve parameter-free vdW-DF descriptions of total energies and structure [74]. Also, AHBR-mRSH delivers QP predictions that employ the same nonlocal-correlation description [93] that is already in wide KS-DFT usage [104, 74, 23]. However, by including some Fock exchange in AHBR-mRSH we also simultaneously correct some density-driven errors, e.g, for transition states, that affects the underlying rev-vdW-DF2 functional [74, 23]. In contrast, the here-introduced KI-CX approach (described below) leaves the GS energies and structure unchanged [28, 29] and therefore tests CX directly in terms of both GS and QP properties.